Could you please help jessica is twice as old as tina. In nine years the sum of their ages will be 66 how old is each today?here's what I have so far... Tina equals t while jessica equals 2t...(T+9)+(2t+9)equals 66
Found 2 solutions by rfer, edjones:
Answer by rfer(10417)

t+9+2t+9=66
3t+18=66
3t=48
t=16
2t=32
your equation works
Good luck
Bob

Answer by edjones(7311)

Let jesica's age=j and tina's=t
j=2t
j+9+t+9=66
2t+9+t+9=66 Substitute 2t for j.
3t+18=66
3t=48
t=16
j=32
.
Ed